OpenStudy (anonymous):

What is the equation of the line that has a slope of -1 and goes through point (0, 5)?

OpenStudy (ranga):

y = mx + b m = -1 y = -x + b passes through (0,5) put x = 0, y = 5 and solve for b

OpenStudy (lalaly):

equation of a line is in the form y=mx+c where m is the slope and c is the y intercept so here slope is -1 so m=-1 so the equation becomes y=-x+c to find c, you substitute (0,5) in x and y 5=-1*0+c so c=5